An object of mass 80kg is whirled round in a vertical circle of radius 2m with a constant velocity of 6m/s. Calculate the maximum and minimum tensions in the string.


1
Expert's answer
2021-02-15T00:58:55-0500
"T_{max}=\dfrac{mv^2}{r}+mg,""T_{max}=\dfrac{80\ kg\cdot(6\ \dfrac{m}{s})^2}{2\ m}+80\ kg\cdot9.8\ \dfrac{m}{s^2}=2224\ N,""T_{min}=\dfrac{mv^2}{r}-mg,""T_{min}=\dfrac{80\ kg\cdot(6\ \dfrac{m}{s})^2}{2\ m}-80\ kg\cdot9.8\ \dfrac{m}{s^2}=656\ N."